OpenStudy (anonymous):

My dad is helping me figure it out and we have solved it- we're just checking to see if we are correct or not! Use elimination to solve the system of equations. {2/3x-3y=1/5 {2x-9y=4

OpenStudy (anonymous):

He thinks that the lines are parallel and have no solution. Are we correct?

OpenStudy (anonymous):

\[2x-9y=4 \]

OpenStudy (anonymous):

\[2x=4+9y \]

OpenStudy (anonymous):

\[x=4/2 + 9/2y \]

OpenStudy (anonymous):

\[ x=2+9/2y\] \[2/3(2+9/2y)-3y =1/5\] \[4/3+18/6y-3y=1/5\] \[4/3+3y-3y =1/5 \] 4/3=1/5 So no solution possible! Am I correct? @Directrix @whpalmer4

Directrix (directrix):

I agree. The lines are parallel; they have the same slope. So, there is no intersection and no common solution.

OpenStudy (anonymous):

Thank you! Hey, I'm starting to get the hang of this. =)

Directrix (directrix):

Glad to hear the good news. Practice is the key.

OpenStudy (anonymous):

=) Yupp!

Directrix (directrix):

Here are some study aids.

OpenStudy (anonymous):

Wow thanks! You're awesome.

Directrix (directrix):

Thanks.

OpenStudy (whpalmer4):

Yes, they are parallel lines, so no intersection and no common solution. Graph attached for your viewing pleasure.